|5x + 9| + 6 = 9, i got one answer but not the other. how do i solve the other answer (my nswer is -6/5)
Answer by lwsshak3(11628)

|5x + 9| + 6 = 9, i got one answer but not the other. how do i solve the other answer (my nswer is -6/5)
**
|5x+9|+6=9
|5x+9|=3
for (5x+9)>0
5x+9=3
5x=-6
x=-6/5
..
for (5x+9)<0
-|5x+9|=3
-5x-9=3
-5x=12
x=-12/5